Operations Report Oct 18th – 20th

MDRS Operations Report 18-20-OCT-2021

Name of person filing report: Shannon Rupert

Reason for Report: Routine

Non-nominal systems: Nothing to report

Action taken for non-nominal systems: Nothing to report

Generator: Nothing to report

ScienceDome Dual Split: Off

Solar—Nominal, supplying all power

Solar— SOC % Last 24 hours (Wednesday):

Average 88

Minimum 73

Maximum 100

Notes on solar: Nothing to report

Propane Reading, station tank – 61 %

Propane Reading, director tank— 70 %

Propane Reading, intern tank— 72 %

Propane Reading, generator— 80 %

Ethanol Free Gasoline – 1 gallons

Water (loft tank): 40 gallons

Water Meter: not checked units

Water (static tank) – 150 gallons

Static to Loft Pump used – no

Water in GreenHab – 10 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 150 gallons

Hab toilet tank emptied: no

Perseverance rover used: yes

Hours: 220.6

Beginning charge: 100

Ending charge: 97

Currently charging: yes

Sojourner rover used: yes

Hours: not noted

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Spirit rover used: no

Hours: 141.0

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Opportunity rover used: no

Hours: 73.0

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: no

Hours: 148.0

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 350.3): none

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? No

CrewCar used and why, where? Yes, to town

Luna used and why, where? Yes, to town

General notes and comments: Nothing to report

Summary of internet: All accounts are nominal.

Summary of suits and radios: Nothing to report

Campus wide inspection, if action taken, what and why? I went to repair what I thought were a few areas where the wire holding the cattle panels needed replacing and I discovered that much of the wire has rusted and broken off and so much of the tunnels need replacement wire. I did some today.

Summary of general operations: The new sign seems to be working better than anything else we have tried, as people have mostly parked and viewed the station from the sign as instructed.

Summary of Hab operations: Nothing to report

Summary of Outpost operations: Got the rest of the work party trash materials on the trailer except for the things that were too big for me to handle alone.

Summary of GreenHab operations: Shade cloth needs to be removed. The plants are protesting phototactically. Spent some time on a new crop project.

Summary of ScienceDome operations: Working on some NASA Spaceward Bound Utah geology projects.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: I went into the Observatory and pressed a button today at Peter’s request. Finally made it to the inner sanctum.

Summary of health and safety issues: Hand is fine. Not healing but also not infected.

Questions, concerns, supplies needed and requests: Mouse is gone.